OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? -2x + 5y = -12 -6x + _ y = _

OpenStudy (anonymous):

@KitKat16 @Kit_Kat_Nat_<3 @Directrix @dan815 @Destinycholie @Donblue22 @Dqswag @waterineyes @HaileyD @birvin73 @Secret-Ninja @SolomonZelman @Zarkon @Mashy @marylou004 @leonardo0430

OpenStudy (secret-ninja):

I'm guessing the underscores are whats missing?

OpenStudy (anonymous):

yup

OpenStudy (secret-ninja):

Well, int the second one, the y is missing so you have to solve for x in the first then put them together to solve for the y and get the solution in the second if I'm not mistaken.

OpenStudy (secret-ninja):

Do you know how to solve for x?

OpenStudy (anonymous):

@Secret-Ninja there is not enough details to solve for x..

OpenStudy (anonymous):

It is given that : the lines are coincident... we must use that..

OpenStudy (secret-ninja):

Well, I know that, but to set it up so that is is on its own side is what I meant.

OpenStudy (anonymous):

well that's all of the question

OpenStudy (anonymous):

For two equations like : \(a_1x + b_1y = c_1\) \(a_2x + b_2y = c_2\) If lines are coincident, then: \[\frac{a_1}{a_2} = \frac{b_1}{b_2} = \frac{c_1}{c_2}\]

OpenStudy (anonymous):

it has to be solved by putting in the rest of the equation that makes it coincident you have to put it in standard form

OpenStudy (anonymous):

So: \[\frac{-2}{-6} = \frac{5}{b_2} = \frac{-12}{c_2}\]

OpenStudy (anonymous):

there are two equations you just have to finish one

OpenStudy (anonymous):

From here, solve for \(b_2\) and \(c_2\)..

OpenStudy (secret-ninja):

-2x + 5y = -12 move the 5y -2x = -12 - 5y (subtracted because it changes when flipped) now you divide the -2 and x so you get: \[x = \frac{ 5y }{ 2 } + 6\]

OpenStudy (anonymous):

so its 5y=-2x+6?

OpenStudy (anonymous):

Second equation must be a multiple of first..

OpenStudy (anonymous):

As you can see, second equation is 3 times first, by looking at \(x\) coefficient.. So, multiply every coefficient of first equation with \(3\), you will get what you require here.

Directrix (directrix):

-2x + 5y = -12 -6x + _ y = _ --------------- If the lines are to be coincident, then they will coincide, be the same line. So, the equations have to be equivalent. Looking at the x terms, 3 times -2x = -6x So, multiply 3 times 5 to get 15 for the missing coefficient of y. Multiply -12 by 3 to get the missing constant term. @help-_-please